ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 241, 2021

Antolín, Y. and L. Paris Transverse properties of parabolic subgroups of Garside groups

501

Avilés, A. See F. Cabello-Sánchez Balodi, M., A. Banerjee and S. Ray Entwined modules over linear

categories and Galois extensions

623

Banerjee, A. See M. Balodi Banerjee, D. and C. Chaudhuri Semi-stable models of modular

curves X_{0}(p^{2}) and some arithmetic applications

583

Barthel, T., G. Heuts and L. Meier A Whitehead theorem for periodic homotopy groups

1

Bauer, T. and B. Greenfeld Prime monomial ideals of algebras of subsemigroups of free nilpotent groups

527

Bienvenu, L., N. Greenberg and B. Monin Bad oracles in higher computability and randomness

229

Blagoveshchenskaya, E. and L. Strüngmann Near-isomorphism for bounded completely decomposable groups

277

Borodulin-Nadzieja, P. See F. Cabello-Sánchez Burns, D. and S. Seo On circular distributions and a conjecture of

Coleman

343

Cabello-Sánchez, F., A. Avilés, P. Borodulin-Nadzieja, D. Chodounský and O. Guzmán Splitting chains, tunnels and twisted sums

955

Cappelletti-Montano, B., A. De Nicola, J. C. Marrero and I. Yudin Almost formality of quasi-Sasakian and Vaisman manifolds with applications to nilmanifolds

37

Chan, W. and S. Jackson The destruction of the axiom of determinacy by forcings on ℝ when Θ Is regular

119

Chaudhuri, C. See D. Banerjee Chodounský, D. See F. Cabello-Sánchez

Christensen, J. D. and E. Wu Smooth classifying spaces  911
Costanzo. D. G. and M. L. Lewis Central Camina pairs  991

De Nicola, A. See B. Cappelletti-Montano Di Vincenzo, O. M., V. R. T. da Silva and E. Spinelli Minimal

varieties of PI-superalgebras with graded involution

869

Diniz, D. See C. Fidelis Ernst, R., C. Esser and Q. Menet U-frequent hypercyclicity no-

tions and related weighted densities

817

Esser, C. See R. Ernst Femić, B. Turaev bicategories and generalized Yetter–Drinfel‘d

modules in 2-categories

395

Fidelis, C., D. Diniz and F. L. de Sousa Identities for the special linear Lie algebra with the Pauli and Cartan gradings

187

Furtado, M. F. and J. P. P. da Silva Positive solution for an indefinite fourth-order nonlocal problem

775

Genevois, A. Algebraic characterisation of relatively hyperbolic special groups

301

Glutsyuk, A., I. Izmestiev and S. Tabachnikov Four equivalent properties of integrable billiards

693

Greenberg, N. See L. Bienvenu Greenfeld, B. See T. Bauer Guzmán, O. See F. Cabello-Sánchez

Haykazyan, L. and J. Kirby Existentially closed exponential fields  89

Heuts, G. See T. Barthel Holmsen, A. F. and D. Lee Radon numbers and the fractional Helly

theorem

433

Hoshi, A. and A. Yamasaki Rationality problem for norm one tori  849
Huang, T. L^{2} vanishing theorem on some Kähler manifolds  147

Izmestiev, I. See A. Glutsyuk Jackson, S. See W. Chan Ji, C. and V. D. Rădulescu Concentration phenomena for nonlinear

magnetic Schrödinger equations with critical growth

465

Jojić, D., G. Panina and R. Živaljević A Tverberg type theorem for collectively unavoidable complexes

17

Karasev, R. Mahler’s conjecture for some hyperplane sections  795

Kirby, J. See L. Haykazyan Lee, D. See A. F. Holmsen Lewis, M. L. See D. G. Costanzo Liebeck, M. W., A. Shalev and P. H. Tiep On the diameters of

McKay graphs for finite simple groups

449

Magee, M. and D. Puder Surface words are determined by word measures on groups

749

Marrero, J. C. See B. Cappelletti-Montano Meier, L. See T. Barthel Menet, Q. See R. Ernst Monin, B. See L. Bienvenu Panina, G. See D. Jojić Paris, L. See Y. Antolín Pène, F. and D. Thomine Central limit theorems for the ℤ^{2}-periodic

Lorentz gas

539

Puder, D. See M. Magee Rădulescu, V. D. See C. Ji Ray, S. See M. Balodi Seo, S. See D. Burns Shalev, A. See M. W. Liebeck da Silva, J. P. P. See M. F. Furtado da Silva, V. R. T. See O. M. Di Vincenzo Simon, R., S. Spież and H. Toruńczyk Games of incomplete

information and myopic equilibria

721

de Sousa, F. L. See C. Fidelis Spież, S. See R. Simon Spinelli, E. See O. M. Di Vincenzo Strüngmann, L. See E. Blagoveshchenskaya Tabachnikov, S. See A. Glutsyuk

Thakur, M. The cyclicity problem for Albert algebras  139

Thomine, D. See F. Pène Tiep, P. H. See M. W. Liebeck Toruńczyk, H. See R. Simon Wu, E. See J. D. Christensen Yamasaki, A. See A. Hoshi Yudin, I. See B. Cappelletti-Montano Živaljević, R. See D. Jojić